# VIPeR: Visual Incremental Place Recognition with Adaptive Mining and Lifelong Learning

### v1

@@ III-B Brain-inspired Memory Bank
To alleviate the notorious catastrophic forgetting in lifelong learning, we first investigate the rehearsal-based approach to refresh the model’s memory about previous environments. Specifically, we present a brain-inspired memory bank $\mathcal{B}=(M,R)$ with a fixed-size image storage space $M$ and a corresponding adjacency matrix $R$. Each element in the adjacency matrix $R_{i,j}$ represents the relative location between place $i$ and place $j$, which is then used to determine whether these two places belong to a positive or a negative pair.
@@
Regarding the fixed-size image storage space $M$, we mimic the human memory system and design the architecture as a composition of a per-environment sensory memory $M^{sn}$, a per-environment working memory $M^{wk}$, and an environment-agnostic long-term memory $M^{lt}$. Our brain-inspired memory bank is capable of improving the model’s adaptability in a new environment while preserving its stability over all previous environments. We illustrate the update process of the proposed brain-inspired memory bank in Fig. 2.
@@
The long-term memory $M^{lt}$, also as a list of size $l^{lt}$, stores the samples of places from all previously visited environments to ensure its environment-agnostic property. When entering a new environment $t$, the long-term memory $M_{t}^{lt}$ is updated following
$M_{t}^{lt}=\omega M_{t-1}^{wk}+\left(1-\omega\right)M_{t-1}^{lt},$ (2)
where $M^{wk}_{t-1}$ and $M^{lt}_{t-1}$ are the working memory and the long-term of environment $t-1$, $\omega$ is a hyper-parameter to control the update ratio.
@@ III-C Relational Memory Aware Synapses
$\Omega^{RMAS}_{t}\approx\frac{1}{N_{t}}\sum_{k=1}^{N_{t}}\left(\frac{\partial\lVert\tilde{S}_{k,t}\rVert_{F}}{\partial\theta}\right)^{2},$ (3)
@@
$L_{RMAS}=\sum_{\theta\in\Theta}^{n}\Omega^{RMAS}_{t-1}\left(\theta_{t}-\theta_{t-1}\right)^{2},$ (4)
@@ III-D Probabilistic Knowledge Distillation
Moreover, we delve into the function regularization approach with knowledge distillation to further enhance the model’s adaptability to unseen environments during lifelong learning. Although previous works like AirLoop [18] and InCloud [19] have also explored knowledge distillation, we believe that their performance is limited due to two reasons. First of all, when selecting memory samples for distillation, these works use images from the current environment. We claim that such a choice is counter-intuitive as the previous model hasn’t seen the current environment, and its performance cannot be guaranteed. Secondly, these works opt for deterministic distillation with angular metrics, which makes the model more focused on local structures of samples.
To select more effective samples, we propose to only use the samples from the long-term memory $M^{lt}$ of our brain-inspired memory bank to perform knowledge distillation. Besides, we present PKD to make the model pay more attention to the global distribution of the samples. To estimate the distribution of the samples in the learned feature space, we first construct the matrices $H_{t}$ and $H_{t-1}$ using the model trained in the current environment $t$ and the frozen model that has been trained on environment $11,2,...,t-$ 1 respectively. Specifically, elements in the matrix are computed as
@@ IV-A Implementation Details
When forming the triplet tuple, we empirically set the number of positive and negative samples to be $m=1$ and $n=5$. And we set the margin in Eq. 1 to be $\delta=1$. Regarding adaptive mining, we prevent the model from always learning with easy samples by explicitly setting the threshold for using easier samples to be higher than the one for using more difficult ones. Especially, we choose $T_{d}=0.02$ and $T_{e}=0.01$.
As for the brain-inspired memory bank, we ensure a fair comparison by setting the total size of our memory bank to be the same as the queue’s length in AirLoop [18]. Particularly, we empirically divide the total size into $l^{sn}=500$ for sensory memory, $l^{wk}=400$ for working memory, and $l^{lt}=100$ for long-term memory. In addition, we set the hyper-parameter in Eq. 2 to be $\omega=0.5$.
Finally, for Eq. 7, we choose equal weights with hyper-parameters $\lambda_{1}=1$ and $\lambda_{2}=1$. We train the VPR methods using the stochastic gradient descent (SGD) optimizer with a learning rate of $0.002$ and a momentum of $0.9$.
Figure 3: Top-$1$ retrieval of the model trained for $T$ environments and evaluated on the first environment. Our VIPeR can better discriminate similar places and exhibits better resilience to catastrophic forgetting.
@@ IV-B Datasets and Evaluation Metrics
Evaluation metrics: In line with AirLoop [18], we use the recall rate at 100% precision as the core metric to evaluate the model’s performance on single environments. To better assess the performance in lifelong learning, we evaluate the model trained after each environment on every environment in the environment sequence $\mathcal{E}$ of length $T$, even the unseen ones. We summarize the results in an evaluation matrix $P\in\mathbb{R}^{T\times T}$ with each of its element $P_{i,j}$ being the recall rate at 100% precision when evaluating the model trained after environment $i$ in environment $j$.
@@ IV-C Evaluation and Discussions
For comparison, we first set a baseline performance with naive finetuning on both of the VPR methods we experimented with. In addition, we also explore the performance of two generic weight regularization methods in lifelong learning, EWC [26] and SI [27]. For more recent methods, we compare our proposed VIPeR to AirLoop [18] and InCloud [19], which are specifically designed for visual incremental place recognition. However, because AirLoop only provided the performance with the combination of VGG-19 [42] and GeM [37], we re-train AirLoop with their official implementation to get its performance with the combination of VGG-19 and NetVLAD [2]. Similarly, since InCloud was only tested with LiDAR point clouds, we replaced its VPR methods with the same one that our VIPeR experimented with and re-trained the InCloud model with their official implementation.
We present the qualitative results in Fig. 3 with VIPeR, AirLoop [18], and InCloud [19] trained for all available environments and evaluated on the very first one to demonstrate the model’s resilience to catastrophic forgetting. Then, we also present quantitative results in all three environments in Table I. It can be seen that naive finetuning and generic weight regularization method EWC [26] and SI [27] exhibit inferior performance, especially in AP and BWT, indicating their inadequate abilities when fighting against catastrophic forgetting in visual incremental place recognition. Furthermore, despite that both AirLoop [18] and InCloud [19] achieved improved performance over baseline methods by utilizing a combination of several lifelong learning methods to mitigate the catastrophic forgetting, they suffer from a notable performance drop when changing from GeM [37] to more discriminative NetVLAD [2].
@@
TABLE I: Performance comparison of lifelong place recognition methods. All the results are obtained with VGG-19 [42] and the ones marked with § are directly taken from AirLoop [18]. boldface and underline indicate the best and second-best results.
Feature Aggregator Method Oxford RobotCar Nordland TartanAir
@@
GeM [37] §Finetune 0.411 -0.066 0.462 0.615 -0.012 0.549 0.754 -0.009 0.730
@@
Figure 4: Performance of different mining strategies on Nordland and TartanAir datasets. “hard2”, “hard5”, “hard10” indicate the performance of using the hard mining strategy with 2, 5, and 10 negative samples.
TABLE II: Comparison of different data mining strategies on the Oxford RobotCar Dataset
@@
random mining 0.454 -0.028 0.481
2 hard mining 0.267 -0.115 0.371
5 hard mining 0.112 -0.010 0.096
10 hard mining 0.161 -0.027 0.161
adaptive mining 0.559 -0.020 0.587
IV-D Ablation Study
We further investigate the contribution of the adaptive mining, brain-inspired memory bank, and PKD loss by replacing the corresponding modules in the base model, i.e. AirLoop [18], one at a time.
Adaptive Mining: To start with, we replace the random mining with hard mining, which is commonly used in VPR tasks, and the proposed adaptive mining. We conduct experiments in all three datasets as above and present the results in Fig. 4 and Table II. The adaptive mining strategy used in our proposed VIPeR exhibits favorable performances in TartanAir [45] and Oxford RobotCar [43], outperforming random mining used in the base model and hard mining used in VPR methods with batched inputs. Although our adaptive mining does not perform as good as hard mining in Nordland [44] when we increasing the number of negative samples, we believe it is mainly caused by the repetitive scenes between different environments in Nordland. We argue that hard mining introduces serious bias into the learning process, which is demonstrated by the drastic performance drop in TartanAir and Oxford RobotCar.
TABLE III: Comparison of different rehearsal methods on the Oxford RobotCar Dataset
@@ IV-D Ablation Study
naive queue 0.454 -0.028 0.481
memory bank 0.494 0.016 0.510
TABLE IV: Comparison of different knowledge distillation methods on the Oxford RobotCar Dataset
@@
Finetune 0.411 -0.066 0.462
RKD 0.454 -0.028 0.481
AKD 0.491 -0.003 0.507
PKD 0.558 0.005 0.567
Brain-inspired Memory Bank: Then, we replace the naive queue used in the base model with the brain-inspired memory bank in the proposed VIPeR and present the quantitative results on Oxford RobotCar [43] dataset in Table  III. It is notable that the brain-inspired memory bank surpasses the naive queue with a large margin across all evaluation metrics. The results demonstrate the superior performance of the brain-inspired memory bank in lifelong learning with good generalizability and resilience to catastrophic forgetting.
Probabilistic Knowledge Distillation: Finally, replace the relational knowledge distillation in the base model with the PKD used in the proposed VIPeR. In addition, we also compare with performance of angular-based knowledge distillation proposed in InCloud [19]. As shown in Table IV, it is evident that our probabilistic distillation largely exceeds the performance of relational knowledge distillation and angular-based knowledge distillation, let alone the naive finetuning strategy. The extraordinary results validate that our distillation approach is capable of grasping the underlying data distribution, which is a big advantage in dealing with catastrophic forgetting and the adaptability when training in unseen environments.
@@ V Conclusions
In this work, we propose VIPeR, a visual incremental place recognition method. We investigate triplet mining strategies suitable for lifelong learning and find a balance between per-environment and cross-environment performance with the proposed adaptive mining. In addition, we mitigate the catastrophic forgetting issues by combining a rehearsal-based approach with a regularization-based approach. In particular, we design a novel brain-inspired memory bank to mimic the human memory system and present a probabilistic knowledge distillation. We conduct extensive experiments on large-scale datasets, the results of which demonstrate the effectiveness of the proposed VIPeR. As for future work, we believe there are many possible directions, like extending the VIPeR model to handle other forms of single-modality input or cross-modality input.

### v3

@@ III-B Multi-stage Memory Bank
To alleviate the notorious catastrophic forgetting in continual learning, we first investigate the rehearsal-based approach to refresh the model’s memory about previous environments. Specifically, we present a multi-stage memory bank $\mathcal{B}=(M,R)$ with a fixed-size image storage space $M$ and a corresponding adjacency matrix $R$. Each element in the adjacency matrix $R_{i,j}$ represents the relative location between place $i$ and place $j$, which is then used to determine whether these two places belong to a positive or a negative pair.
@@
Regarding the fixed-size image storage space $M$, we mimic the human memory system and design the architecture as a composition of a per-environment sensory memory $M^{sn}$, a per-environment working memory $M^{wk}$, and an environment-agnostic long-term memory $M^{lt}$. Our multi-stage memory bank is capable of improving the model’s adaptability in a new environment while preserving its stability over all previous environments. We illustrate the update process of the proposed multi-stage memory bank in Fig. 2.
@@
The long-term memory $M^{lt}$, also as a list of size $l^{lt}$, stores the samples from all previously visited environments to ensure its environment-agnostic property. When entering a new environment $t$, the long-term memory $M_{t}^{lt}$ is updated with $M_{t}^{lt}=\omega M_{t-1}^{wk}+\left(1-\omega\right)M_{t-1}^{lt}$, where $M^{wk}_{t-1}$ and $M^{lt}_{t-1}$ are the working memory and the long-term of environment $t-1$, $\omega$ is a hyper-parameter to control the update ratio.
@@ III-C Relational Memory Aware Synapses
$\Omega^{RMAS}_{t}\approx\frac{1}{N_{t}}\sum_{k=1}^{N_{t}}\left(\frac{\partial\lVert\tilde{S}_{k,t}\rVert_{F}}{\partial\theta}\right)^{2},\vspace{-1ex}$ (2)
@@
$L_{RMAS}=\sum_{\theta\in\Theta}^{n}\Omega^{RMAS}_{t-1}\left(\theta_{t}-\theta_{t-1}\right)^{2},\vspace{-1ex}$ (3)
@@ III-D Probabilistic Knowledge Distillation
Moreover, we delve into the function regularization approach with knowledge distillation to further enhance the model’s adaptability to unseen environments during continual learning. Although previous works have also explored knowledge distillation, we believe that their performance is limited due to two reasons. First of all, when selecting memory samples for distillation, CCL [16] use images from the past environment. We believe that using data unseen by the old model for knowledge distillation helps to extract the common knowledge, thereby enhancing the model’s generalization ability. Secondly, these works like AirLoop [14] and InCloud [15] opt for deterministic distillation with angular metrics, which makes the model more focused on local structures of samples.
To select more effective samples, we propose to use the samples from our multi-stage memory bank $M$ to perform knowledge distillation. Besides, we present PKD to make the model pay more attention to the global distribution of the samples. To estimate the distribution of the samples in the learned feature space, we first construct the matrices $H_{t}$ and $H_{t-1}$ using the model trained in the current environment $t$ and the frozen model that has been trained on environment $11,2,...,t-$ 1 respectively. Specifically, elements in the matrix are computed as
@@ IV-A Implementation Details
When forming the triplet tuple, we empirically set the number of positive and negative samples to be $m=1$ and $n=5$, and we set the margin to be $\delta=1$. Regarding adaptive mining, we prevent the model from always learning with easy samples by explicitly setting the threshold for using easier samples to be higher than the one for using more difficult ones. Especially, we choose $T_{d}=0.05$ and $T_{e}=0.03$.
As for the multi-stage memory bank, we ensure a fair comparison by setting the total size of our memory bank to be the same as the queue’s length in AirLoop [14]. Particularly, we empirically divide the total size into $l^{sn}=500$ for sensory memory, $l^{wk}=400$ for working memory, and $l^{lt}=100$ for long-term memory. In addition, we set the hyper-parameter in the memory bank to be $\omega=0.5$.
Finally, we train the VPR methods using the stochastic gradient descent (SGD) optimizer with a learning rate of $0.002$ and a momentum of $0.9$. All experiments in this paper are conducted on a single NVIDIA RTX 4090 GPU and each dataset was trained for 2 to 3 hours.
@@ IV-B Datasets and Evaluation Metrics
Evaluation metrics: In line with AirLoop [14], we use the recall rate at 100% precision as the core metric to evaluate the model’s performance on single environments. To better assess the performance in continual learning, we evaluate the model trained after each environment on every environment in the environment sequence $\mathcal{E}$ of length $T$, even the unseen ones. We summarize the results in an evaluation matrix $P\in\mathbb{R}^{T\times T}$ with each of its element $P_{i,j}$ being the recall rate at 100% precision when evaluating the model trained after environment $i$ in environment $j$.
@@ IV-C Evaluation and Discussions
We first set a baseline performance with naive finetuning on both of the VPR methods we experimented with. In addition, we also explore the performance of two generic weight regularization methods in continual learning, EWC [23] and SI [24]. For more recent methods, we compare our proposed VIPeR to AirLoop [14] and InCloud [15], which are specifically designed for visual incremental place recognition. Since AirLoop only supports the combination of VGG-19 [50] and GeM [30], and InCloud takes in only LiDAR point clouds, both are modified and re-trained, using their official implementation, to get comparison results. Despite our eager to compare with BioSLAM [18], neither of their code or data were publicly available.
We present the qualitative results in Fig. 3 with VIPeR, AirLoop [14], and InCloud [15] trained for all available environments and evaluated on the very first one to demonstrate the model’s resilience to catastrophic forgetting. Then, we also present quantitative results in all three environments in Table I. It can be seen that naive finetuning and generic weight regularization method EWC [23] and SI [24] exhibit inferior performance, especially in AP and BWT, indicating their inadequate abilities when fighting against catastrophic forgetting in visual incremental place recognition. Furthermore, despite that both AirLoop [14] and InCloud [15] achieved improved performance over baseline methods by utilizing a combination of several continual learning methods to mitigate the catastrophic forgetting, they suffer from a notable performance drop when changing from GeM [30] to more discriminative NetVLAD [2].
@@
We further highlight the importance of continual learning within the context of foundation models, a necessity that becomes evident when examining models like AnyLoc [42], a DINOv2-based place recognition model with broad compatibility across diverse data types. To ensure a fair comparison, we utilize AnyLoc pre-trained specifically for urban and indoor environments and adjust our backbone to DINOv2 [38]. We challenge our VIPeR by conducting experiments not only on single datasets but also across multiple datasets—TartanAir, Nordland, and RobotCar—testing the model’s robustness under varied conditions. Table II presents our quantitative results, where VIPeR outperforms AnyLoc by a substantial margin across all three datasets. Notably, when using the VGG [50] backbone, VIPeR’s cross-dataset training slightly diminishes performance. However, with DINOv2 as the backbone, cross-dataset performance not only matches but surpasses single-dataset results, demonstrating enhanced effectiveness of continual learning in conjunction with visual foundation models for robust place recognition.
@@
Backbone Feature Aggregator Method Oxford RobotCar Nordland TartanAir
[…]
VGG [50] GeM [30] §Finetune 0.411 -0.066 0.462 0.615 -0.012 0.549 0.754 -0.009 0.730
@@
Model Method TartanAir Nordland RobotCar
DINOv2 [38] + NetVLAD [2] AnyLoc-urban [42] 0.678 0.543 0.647
AnyLoc-indoor [42] 0.679 0.483 0.602
@@
VIPeR (M.D.M.S)2 0.807 0.711 0.680
@@
[…]
1 S.D.M.S. stands for single-dataset, multiple-scene, means a model is trained for one dataset with different scenes;
2 M.D.M.S. is multiple-dataset, multiple-scene, means the model is trained across all three datasets in the order of TartanAir, Nordland, and RobotCar.
IV-D Ablation Studies
Figure 4: Performance evaluated in the first environment after training in all environments with different mining strategies, highlighting forgetfulness. “hard5”, “hard10” indicate the performance of using the hard mining strategy with 5, and 10 negative samples.
[…]
Method TartanAir Nordland RobotCar
[…]
VIPeR (random mining) 0.784 0.649 0.547
VIPeR (5 hard mining) 0.786 0.702 0.537
VIPeR (10 hard mining) 0.778 0.725 0.495
VIPeR (smart mining) 0.789 0.681 0.514
VIPeR (naive queue) 0.758 0.632 0.460
VIPeR (RKD) 0.793 0.698 0.520
[…]
Adaptive Mining: We first validate the contribution of our adaptive mining by comparing to the performance using random mining, hard mining, and smart mining [49]. With the results from the three datasets shown in Fig. 4 and Table III, we can see that the adaptive mining strategy proposed in our VIPeR exhibits favorable performances, especially in TartanAir [53] and Oxford RobotCar [51]. Although our adaptive mining does not perform as good as hard mining with 10 negative samples in Nordland [52], we believe it is mainly caused by the less divergent scenes in Nordland, reducing the impact of the bias.
Multi-stage Memory Bank: Then, we replace the multi-stage memory bank with the naive queue used in AirLoop [14] and InCloud [15], and present the quantitative results in Table  III. It is notable that the multi-stage memory bank surpasses the naive queue by a large margin across all datasets, demonstrating the superior performance of the multi-stage memory bank in continual learning with good generalizability and resilience to catastrophic forgetting.
Probabilistic Knowledge Distillation: Finally, the contribution of PKD is evalutated by replacing with relational knowledge distillation (RKD) from AirLoop [14] and angular-based knowledge distillation (AKD) from InCloud [15]. As shown in Table III, it is evident that our PKD exceeds the performance of RKD and AKD, validating that our distillation approach is capable of grasping the underlying data distribution, which is a big advantage in dealing with catastrophic forgetting and the adaptability when training in unseen environments.
@@ V Conclusions
In this work, we present VIPeR, a visual incremental place recognition method. We dive into triplet mining strategies and find a balance between per-environment and cross-environment performance with adaptive mining. In addition, we design a novel multi-stage memory bank and present a probabilistic knowledge distillation to mitigate the catastrophic forgetting. Through extensive experiments on large-scale datasets, we demonstrate the superior of the our VIPeR over not only other visual increment place recognition methods, but also visual foundation model-based methods. As for future work, we believe there are many possible directions, like extending the VIPeR model to handle other forms of single-modality input or cross-modality input.